A shell is fired at an angle of 20 degree to , an above, the horizontal from the top of a cliff 90 meter high , the velocity of the shell is 150 meter per second, calculate the distance from the foot of the cliff where it strikes the ground and 2) calculate the highest point above the ground reach by the shell


1
Expert's answer
2022-03-29T05:33:59-0400

"h=h_0+\frac{v_y^2}{2g}=134~m,"

"t=\frac{v_y}g+\sqrt{\frac{2h}g}=10.5~s,"

"l=v_xt=1360~m."